OpenStudy (anonymous):

write in lowest terms : 28x^6 / 98x^3

OpenStudy (anonymous):

\[\frac{ 28x ^{6} }{ 98x ^{3} }\]

OpenStudy (abb0t):

\[\frac{ a^m }{ a^n }=a^{m-n}\]

OpenStudy (abb0t):

Also, what's the greatest common number between 28 and 98? Numbers of 28: 1, 2, 4, 7, 14, 28 Numbers of 98: 1, 2, 7,....98 Lol. Sorry I can't think of any numbers in between but i'm sure you can figure this out.

OpenStudy (anonymous):

the GCM is 14